OpenStudy (kpop4life123):

The hypotenuse of a right triangle is 12, and one of the legs is 4. Find the length of the missing leg to the nearest hundredth.

OpenStudy (anonymous):

so 12^2=144then 4^2=16 144+16=160^2=25600

OpenStudy (anonymous):

oh wait.......its 13

OpenStudy (anonymous):

bc 160 yroot 2=12.6 then the nearest hundreth is 13

OpenStudy (tanabugg):

YEah I was right

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no no its 12.5

OpenStudy (anonymous):

I am going with 12.5

OpenStudy (anonymous):

u cant round that more

OpenStudy (tanabugg):

or 12.6

OpenStudy (anonymous):

I think 12.5

OpenStudy (kpop4life123):

im so confused

OpenStudy (tanabugg):

What are you choices ?

OpenStudy (kpop4life123):

it doesnt have any choices

OpenStudy (anonymous):

so after all this u get 160 but u divide by 2 so u get 12.5

OpenStudy (anonymous):

www.wikihow.com/Find-the-Length-of-the-Hypotenuse

OpenStudy (tanabugg):

Square root of 160 is 12.64911064067352 so

OpenStudy (tanabugg):

You don't divide when you do the Pythagorean Theorem

OpenStudy (tanabugg):

a^2 + b^2 = c^2 Therefore, 12*12= 144 4*4= 16 144+ 16= 160 Square root of 160 is 12.64911064067352